# Sequential Quantum Measurements and the Instrumental Group Algebra

### v1

@@ I Introduction
I.1 Prelude
I.1.1 There is a need to get past the idea of instantaneous measurement.
In quantum theory, observables and states are distinct objects, each a certain kind of operator over a Hilbert space. In particular circumstances, an observable can correspond to directly to a choice of Hermitian operator and the act of observation can be understood to be the random appearance of one of its eigenvalues with a probability that can be calculated by the Born rule. In this case, the appearance of the eigenvalue is also understood to signify that the state has projected onto the corresponding eigenspace. This idea of observation, in particular where the post-measurement state is projected, is called a von Neumann measurement or the Luders rule [4, 5, 6]. Useful as it is, the von Neumann-Luders rule fails to make sense of perhaps the most fundamental of observables, the position of a dynamic particle, because the eigendistributions of the position operator are not square-integrable and therefore not sensible state-preparations. (For a more detailed explanation, see the end of section II.1.)
I.1.2 With continuous measurements, what kinds of instruments are possible?
with
where $T$ is the overall time duration of the measurement, $\vec{L}\equiv(L_{1},L_{2},\ldots,L_{n})$ is a finite tuple of Lindblad operators (which generalize the Hermitian observable), $d\vec{W}_{[0,T)}\equiv\{d\vec{W}_{t}\}_{t=0}^{T-dt}$ is a sequence of Wiener increments (often called the measurement record) representing the raw data output by the instrument, $\mathcal{D}\mu[d\vec{W}_{[0,T)}]$ is the Wiener measure, and $\hat{X}$ denotes a representation of the abstract Lie algebra element $X$. (See Appendix A for a derivation of $\hat{\updelta}^{(L)}$.)
where
I.1.3 There is an analytic structure to applying instruments in sequence.
where $d_{\mathrm{L}}x$ is the left-invariant Haar measure, $\delta_{\text{IG}}(x)$ is the IG delta-function, and the dependence on any Lindblad operators has been dropped ($\gamma=\gamma^{\hat{L}}$). (See Appendix C for details on the delta-function.) What the KOD essentially does is count the number of measurement records that give the same Kraus operator $x$ at any given time $T$, therefore collecting the actions of a continuous measuring instrument.
$\mathcal{I}_{T}=\Big\{d_{L}x\,D_{T}(x)\mathcal{O}_{x}\Big\}_{x\in\text{IG}}$ (6)
I.1.4 There is a need to get past the state-centric approach to continuous measurement.
        Comparison of the KOD FPKE and the Lindblad master equation
IV) KODs are in the IGA Density operators are in the dual of the von Neumann algebra
※(§IV.4) Fokker-Planck-Kolmogorov equation Lindblad master equation
※(§IV.4) Generator for a diffusive process
IV.4) Generator for a jump process
IV.4) Kolmogorov-Lindblad intertwining relation
III.3) Instrument element irreps of the IG
or equivalently
is indeed a solution to equation 12. What then remains is the use of a few group-analytic tools which are explained in section IV, particularly section IV.4. The steps involve taking a derivative of the total operation with respect to time, applying the KOD FPKE, using the left-translation invariance of the measure ($d_{\mathrm{L}}(e^{hX}x)=d_{\mathrm{L}}x$) to move the right-invariant derivatives from the KOD to the superoperators $\mathcal{O}_{x}$, and finally using the Kraus-operator representation property ($K_{yx}=K_{y}K_{x}$) to observe the intertwining relations,
$\underleftarrow{L}[K_{x}]=\hat{L}K_{x}$ (15)
and therefore
I.1.5 The structure of sequential measurement is the IG convolution.
$\mathcal{Z}_{t}=e^{\mathcal{D}\kappa t}$ (17)
and has the well-known (one-parameter) group property
$\mathcal{Z}_{t+\Delta t}=\mathcal{Z}_{\Delta t}\circ\mathcal{Z}_{t}.$ (18)
where * denotes the IG convolution,
where $d_{\mathbb{R}}x$ is the right-invariant Haar measure of the IG. In terms of the right-invariant measure, this convolution can be recognized as the Chapman-Kolmogorov equation of a right-invariant Markov process. In terms of the left-invariant measure, this convolution allows the right-invariant Markov process to be converted into simple operator calculus, as will be explained in just a moment. Right-invariance is the equivalent (in IG language) to the property that the measuring process is nonadaptive, i.e. invariant to the Kraus operator that just preceeds the next step in time.
$\mathcal{Z}_{D}\equiv\int_{\text{IG}}d_{\mathrm{L}}y\,D(y)\mathcal{O}_{y},$ (21)
then for any second instrument with KOD $G(x)$
$\mathcal{Z}_{G}\circ\mathcal{Z}_{D}=\mathcal{Z}_{G*D}$ (22)
$\mathfrak{Z}^{\text{\Ankh}}_{t}=e^{\mathfrak{D}^{\text{\Ankh}}\kappa t}$ (25)
with the one-parameter group property
$\mathfrak{Z}^{\mathrm{L}}_{G}[D](x)\equiv G*D(x).$ (27)
$\mathfrak{L}_{g}[D](x)\equiv D(g^{\,\text{-}\hskip-1.0pt1}x).$ (29)
Summary of the Various Group Properties
From KODs to Ultraoperators From States to Superoperators
Invariant Chapman-Kolmogorov Equation
$D_{\Delta t}*D_{t}=D_{t+\Delta t}$$\Big|$
Instrumental Group Algebra
I.1.6 The home of the KOD is the IG algebra.
$G||_{1}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,|G(x)|,$ (32)
The superoperators $\mathcal{Z}_{f}$ defined in equation 21 and the hyperoperators $\mathfrak{Z}^{\mathrm{L}}_{f}$ defined in equation 30 are both representations of the IGA (by virtue of linearity and equations 22 and 28). However, between the two it is only the superoperators $\mathcal{Z}_{f}$ that compose a $\ddagger$}-representation and only the ultraoperators $\mathfrak{Z}^{\mathrm{L}}_{f}$ that compose an -representation. The representations of these involutions are the adjoints of the Hilbert-Schmidt and Kolmogorov inner products
[…]
There is apparently a canonical way to derive a C*-algebra in this case, the so-called enveloping C*-algebra [30, p. 47]. The exact meaning and correspondence of the enveloping $\text{C}^{\ddagger}$-algebra and the enveloping $\text{C}^{\text{\Ankh}}$-algebra require further investigations that are already beyond the purpose of the current paper.
IG convolutionIV.1)
IV.3) IG inner product: $(g,f)_{\text{IG}}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,\,\overline{g(x)}f(x)$ $(B|A)_{\mathcal{H}_{\lambda}}\equiv\operatorname{tr}_{\mathcal{H}_{\lambda}}B^{\dagger}A$ :Hilbert-Schmidt inner product
IV.3) Kolmogorov adjoint: $\big(g,\mathfrak{Z}^{\text{\Ankh}}[f]\big)_{\text{IG}}\equiv\big(\mathfrak{Z}[g],f\big)_{\text{IG}}$ $\big(B\big|\mathcal{Z}^{\ddagger}(A)\big)_{\mathcal{H}_{\lambda}}\equiv\big(\mathcal{Z}(B)\big|A\big)_{\mathcal{H}_{\lambda}}$ :Hilbert-Schmidt adjoint
IV.3) Gelfand involution: $f^{\text{\Ankh}}(x)\equiv\overline{f(x^{\,\text{-}\hskip-1.0pt1})}$ $f^{\ddagger}(x)\equiv\overline{f(x^{\dagger})}$ :Cartan involution (§IV.3)※
IV.2) left-translation ultraoperator: $\mathfrak{L}_{a}[f](x)\equiv f(a^{\,\text{-}\hskip-1.0pt1}x)$ $\mathcal{O}_{a}\equiv K_{a}\odot K_{a}^{\dagger}$ :CP superoperator (§III.3)※
IV.2) left-convolution ultraoperator: $\mathfrak{Z}^{\text{\Ankh}}_{f}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,\,f(x)\,\mathfrak{L}_{x}$ $\mathcal{Z}_{f}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,\,\overline{f(x)}\,\mathcal{O}_{x}$ :IGA superoperator (§IV.3)※
I.2 The Commutative Analog
@@
The Instrumental Group (IG), Kraus-Operator Density (KOD), IG convolution, IG Algebra (IGA), convolution ultraoperators, and their relationship to the usual density operators and channel superoperators of standard quantum theory are a lot to swallow for those who have yet to ask for an autonomous, general theory of sequential measurement. If, however, we consider a commutative IG, then all of these tools and concepts can easily be recognized by anyone familiar with classical real analysis. To see the forest from the trees, namely the universal measuring process from the quantum channels, it therefore seems quite instructive to go through this. The baroque notation and strange factors of 2 will look funny, but they are chosen with the intention of making the analogy with the noncommutative case as obvious as possible. It is very interesting to note how almost all of these key elements seem to stem from Laplace.
@@
in exact analogy with equations 9 and 25. Meanwhile, the Kraus operators and instrument elements (equation 55) would also be denoted

### v2

@@ I Introduction: The Instrument Manifold Program
The Instrument Manifold Program (IMP) is a method for analyzing the measuring instruments that are generated by continuous measurement, primarily of the diffusive type. This method originally emerged from the discovery that the measurement of spin-direction corresponding to the spin-coherent positive-operator-valued measure (POVM) [4, 5] can be realized universally (that is for any, arbitrarily large, total angular momentum quantum number ‘$j$’) by a diffusive simultaneous measurement of the three orthogonal components of total angular momentum [1, 6, 7], now called the isotropic spin measurement (ISM) [1, 3]. Such a measurement of the spin-direction is a form of informationally-complete quantum tomography analogous to the measurement of phase-point for continuous-variable systems such as a single mode of light, where the phase space for spin is a sphere instead of a plane. On the measurement of phase-point as a form of informationally-complete quantum tomography, there have been several approaches [8, 9, 10]: the Arthurs-Kelly measurement [11, 12], optical heterodyne [13], optical homodyne tomography [14], and diffusive optical heterodyne [15, 16]. While [4] showed that the spin-analog of [11] does not work for $j>1/2$, the closest phase-point analog to ISM of the approaches just mentioned would be diffusive optical heterodyne [15, 16].
The exact phase-point analog of ISM seems to have first been examined in [17], although the authors did not understand that this measurement resulted in the standard coherent POVM. Therefore [2] was written to demonstrate how the IMP is able to analyze extensively the behavior and result of this diffusive measurement, now called SPQM (the simultaneous momentum and position or “P and Q” measurement) in honor of Antoine Barchielli at the University of Milan, whom the author considers the pioneer of the theory of diffusive measuring instruments [17, 18, 19]. In the application to SPQM [2], many of the key elements of the IMP were much easier to give explicit expressions than in ISM [1], because the canonical commutation relations are simpler than the angular-momentum commutation relations.
Having explored ISM and SPQM extensively, [3] was written to explain how the IMP is able to analyze the general problem of simultaneously measuring non-commuting observables, what in this paper will be refered to as diffusively measuring non-commuting observables simultaneously (DMNCOS) to contrast it from the other types of simultaneous measurements of noncommuting observables mentioned [11, 13, 12]. Fundamentally, the general problem of DMNCOS boils down to whether the time-ordered exponential that maps any particular continuous measurement record to its instrument element can be computed. Geometrically, this is related to whether the manifold of instrument elements is finite-dimensional or not. If it is, the instrument was said to be principal, else it was deemed chaotic [3].
Mathematically, the IMP consists of one main concept, two basic tools, and then a large handful of technical details. The main concept is what the author calls the instrumental group ($IG$) and the basic tools are the so-called Kraus-operator density (KOD) and the modified Maurer-Cartan stochastic differential (MMCSD). The IG is a Lie group that can be identified to contain the elements of the measuring instrument at all times. The KOD is a distribution function of the IG with respect to its (left-invariant) Haar measure and describes at every moment in time how the Kraus operators (or more general elements of the instrument) are distributed. The MMCSD is a tool that is particularly well-suited for handling the stochastic differential equations that describe Brownian motion within a group, allowing one to compute/locate within the IG the instrument element that results from any given continuous measurement record.
An honorable mention must be given to a work by Prahlad Warszawski, et al. [20]. Although the authors were not able to uncover the calculus and analysis native to the IG, they certainly recognized its existence for continuous variable systems. With this recognition, they had some key insights into the nature of the KOD (or the closely related ostensible distribution as they call it) and how to calculate the POVM of a continuous measurement. The author of this paper is incredibly relieved to find that there is at least one set of authors out there who has begun to recognize the significance of the IG.
Returning to the IMP, the tools and perspectives that come with it still require almost anyone who works in quantum measurement to step out of their comfort zone because these basic tools are both non-standard and rather sophisticated. The purpose of this paper is to focus on the KOD and explain it from a different perspective, centering the discussion on the more discrete concept of sequential measurement rather than on continuous measurement as considered previously. By doing this, much of the basic structure of combining instruments in sequence (either discretely or continuously) can be addressed more directly. By stepping away from continuous processes and into sequential processes, the tools of stochastic calculus are replaced by the tools of functional analysis. Specifically, stochastic differential equations and Kolmogorov equations [3] are replaced by the multiplication of IG elements and the group-convolution of functions on the IG.
Wonderfully, every Lie group defines a form of analysis that generalizes standard real and complex analysis. If real analysis could be boiled down to only five essential tools, arguably they are the derivative, the integral, the Taylor series, the Fourier expansion, and the convolution. All five of these tools obtain their structure from one essential property of the real line, $\mathbb{R}$, that it forms a group under addition, thus giving rise to the notion of translation, $x\mapsto x+a$. The first four are perhaps the most fundamental: The usual derivative $d/dx$ is the unique translation-invariant vector field, the integral $\int dx$ is the unique translation-invariant (Haar) measure, the exponential operator $e^{a\frac{d}{dx}}$ of the Taylor series is the translation generated by $d/dx$ for variable $a$, and the exponential kernel $e^{ikx}$ of the Fourier transform as a function of $x\in\mathbb{R}$ is an irreducible representation of $\mathbb{R}$ for every $k$. The convolution is then the crown within which these fundamental jewels sit, combining and intertwining in glorious ways. All of these tools still exist if the real line is replaced by a general finite-dimensional Lie group and this paper demonstrates how these tools of general group analysis are directly relevant to the problem of sequential measurements, both discrete and continuous.
The flow of this paper is therefore as follows:
For those in the earlier stages of research in theoretical quantum science, a quick summary of all the more standard tools in quantum science that will be needed — instruments, channels, and Lindbladians — can be found in appendix A. Having this standard theory under one’s belt, the following sections move into the newer concepts of the IMP [3], where in this paper these newer concepts will be framed in terms of group analysis and held together through the central concept of group convolution, the analytic rule for combining instruments in sequence. This machinery of the IMP is relatively heavy, so sections II and III essentially act as a second and third introductions with increasing mathematical intensity.
Section II revisits the much older, classical problem of invariant heat diffusion. However, with a modern perspective and notation, the approach should make clear that exactly the same structures found in real analysis (though noncommutative in general) are present in the general problem of continuous and sequential measuring instruments.
Section III then begins to march into the general theory of continuous and sequential measurement, albeit in a streamlined fashion, where the focus is to introduce all of the details, while not digging into them just yet for the sake of displaying how all of the key ideas follow from one another. Section III.1 starts off with a discussion of instantaneous measurements and their limitations. Section III.2 then moves into the general problem of continuous measurement in terms of the basic time-ordered exponential that maps continuous measurement records to their corresponding instrument elements or Kraus operators, emphasizing the fundamental concept of the instrumental group (IG) and its universal nature. Section III.3 then introduces the Kraus-operator density (KOD) and the Kolmogorov equation that governs its evolution in time when the measurements are continuous. Section III.4 then begins to explain how the more standard decoherence channels and their Lindblad master equations are in fact representations of the KOD and its Kolmogorov equation. Section III.5 then begins to move the discussion away from continuous measurements and their differential equations to sequential measurements and their group convolutions, in particular explaining that the composition of decoherence channels is a representation of the group convolution of KODs. Finally, section III.6 explains how the group convolution promotes the space of absolutely-integrable functions of the IG (which include the KOD) to a group algebra, the so-called instrumental group algebra (IGA).
Having given the reader a first exposure to all the technical elements involved in completely recasting the analysis of continuous measurement in terms of the IG and KOD, we then start over from the beginning, with the intention of explaining the details.
With the dimension of time, measurements can be put together in a variety of ways — continuously, simultaneously, and sequentially — which can be considered to take place in an instrumental group (IG), the subject of section IV:
Section IV.1 explains the fundamental problem of continuous measurements, calculating the time-ordered product of a sequence of weak Kraus operators. It begins by immediately considering a sequence of two instruments as the convolution of their operation sets. The two weak instruments that generate the most common types of continuous Markov process are introduced, the jumpy Poisson process and the diffusive Wiener process. (Appendix B includes a detailed derivation of these two weak instruments from basic system-meter interactions.) The IG is then defined and the relation between the free product of groups and the convolution of operation sets is explained. Then the fundamental problem is stated more precisely and the topics of automation and chaos are briefly discussed.
Section IV.2 explains the basic observation that the Kraus operators of weak measurements commute, even when the Lindblad operators that generate them do not. With this observation in mind, it then makes perfect sense to consider multiple weak measurements with noncommuting Lindblad operators as simultaneous. Such simultaneous measurements can be run for finite amounts of time, thus generating new types of instruments such as those generated by the general “DMNCOS” procedure. Critically, it is then explained that the finite-time instrument that is defined by such a procedure has its definition solely in the commutation relations of the Lindblad operators that generate their Kraus operators, therefore making the structure of the instrument universal, i.e. independent of Hilbert space. This emphasis is very important because it reinforces the basic fact that in physical examples, these commutation relations are based in physical observables that are more fundamental than the Hilbert spaces on which any of these Kraus operators act, a point that is echoes back to a point made by Kraus, referring to “ […] Bohr, who repeatedly stressed the importance of the classical nature of measuring instruments for the understanding of quantum mechanics.” [21, p. 2] A brief mention is then made to the IG for simultaneously measuring a canonical pair of phase-space quadratures and the IG for simultaneously measuring the three orthogonal angular momentum components and the reader is referred to previous work for further details.
Section IV.3 then finishes with a fresh start, assuming an IG from the get-go and considering the instrument as a representation of the IG. This introduces the notion of a Kraus-operator distribution, a precursor to the Kraus-operator density that does not yet have the properties of a function. The convolution of two distributions over the IG is then defined and related to the operation-set convolution of two instruments in sequence. Finally, it is pointed out that this convolution of distributions already expands what it means for the total operations of a sequence of instruments to satisfy a group property.
Just like the real line, the IG is the domain of an entire suite of analytic tools — with invariant integrals, invariant derivatives, function spaces, and intertwined differential-algebraic equations — where functions such as the Kraus-operator density (KOD), the primary object of section V, can be calculated, propogated, and varied:
Section V.1 begins by introducing the all-important Haar integral measures. These measures allow one in turn to define the convolution of two functions and the density function of a distribution. The relationship between the convolution of functions and the convolution of distributions is explained. Finally, the definition of the Kraus-operator density (KOD) is fixed relative to the left-invariant measure, a choice that stems from the convention that time is considered to move from right to left when multiplying a sequence of Kraus operators.
Section V.2 introduces the left- and right-translation ultraoperators. With them, convolution by a function can also be considered as an ultraoperator. The relations between these left- and right-convolution ultraoperators amongst themselves and each other are discussed. Then right- and left-invariant derivatives are defined. It is explained how these invariant derivatives generate the translation ultraoperators in the same analytic way that the usual derivative of a real line generates the Taylor expansion. Finally, ultraoperator commutators of the invariant derivatives are discussed in terms of the Lie algebra of the IG.
Section V.3 discusses the function spaces that are defined by the IG and its left-invariant measure. Among the various norms one could give the IG-functional space, it is only the $\ell^{1}$ norm that is closed under convolution and the IG algebra (IGA) is finally defined as the Banach algebra of absolutely-integrable functions. The IG inner product is then introduced, the Kolmogorov adjoint is defined, and the Kolmogorov adjoints of the various ultraoperators are considered. Then the IGA is shown to have two involutions, named after Cartan and Gelfand, which respect the convolution product if the IG is unimodular. Finally, their relations to the Kolmogorov ultraoperator and Hilbert-Schmidt superoperator adjoints are discussed.
At very last, section V.4 returns to the problem of continuous measurement and uses the ultraoperator formalism to rederive the KOD Kolmogorov equations in a style that is much like the reverse of section II. The various manifestations of the Markovian group property — either by composition, convolution, or addition — are summarized. The role of the instrument elements in intertwining the differential ultraoperators with the algebraic superoperators is then introduced. This fundamental intertwining relation is then used to formally derive the KOD Kolmogorov forward equation, showing along the way that the instrument elements also intertwine the Markov ultraoperator with the quantum channel superoperator. The Kolmogorov generator is then calculated for both diffusive and jump processes. Finally, it is shown that the instrument elements also intertwine both Kolmogorov generators with the Lindbladian dissipator, as one might expect, thus transforming the differential Kolmogorov equation into the algebraic Lindblad master equations, in noncommutative analogy with the Laplace transform.
The paper then finishes with some concluding remarks in section VI.
II A Commutative Analogy: The Real Line as an Instrument
@@
The Instrumental Group (IG), Kraus-Operator Density (KOD), IG convolution, IG Algebra (IGA), and convolution ultraoperators to come, as well as their relationship to the usual density operators and channel superoperators of standard quantum theory, are a lot to swallow for those who have yet to ask for an autonomous, general theory of sequential measurement. If, however, we consider a commutative IG, then all of these tools and concepts can easily be recognized by anyone familiar with classical real analysis. To see the forest from the trees, namely the universal measuring process from the quantum channels, it therefore seems quite instructive to go through this. The baroque notation and strange factors of 2 will look funny, but they are chosen with the intention of making the analogy with the noncommutative case and its application to sequential measurement as obvious as possible. It is very interesting to note how almost all of these key elements seem to stem from Laplace.
@@
in exact analogy with equations 32 and 48. Meanwhile, the Kraus operators and instrument elements (equation 17) would also be denoted